THE VISBLE SPECTRUM OF TITANIUM DIOXIDE

XIUJUAN ZHUANG AND \underline{TIMOTHY C. STEIMLE} , Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287; RAMYA NAGARAJANN AND JOHN P. MAIER , Department of Chemistry, University of Basel, Basel Switzerland.

Bulk TiO_{2} is a widely used photo-activated catalytic material, yet poorly understood. Much of the motivation for studies of molecular TiO_{2} is the observation^{a} that there is a smooth correlation of the molecular electronic states to the band gap of the bulk. The field-free energy levels of the ground state of the monomer have been fully characterized by microwave spectroscopy.^{b} Here we report on the visible spectrum in the region between 18200 cm^{−1} to 18750 cm^{−1} of a cold molecular beam sample of TiO_{2} using laser induced fluorescence detection and mass-selected REMPI. Bands at 18240 cm^{−1}, 18411 cm^{−1} and 18470 cm^{−1} were recorded at a resolution of 40 MHz and rotationally analyzed. The dispersed fluorescence of 18411 cm^{−1} and 18470 cm^{−1} bands were analyzed to produce a set of vibrational parameters for the ground state. The optical Stark spectra of the 18411 cm^{−1} and 18470 cm^{−1} bands were recorded and analyzed to determine permanent electric dipole moments and compared with the results for the band at 18655 cm^{−1}^{c}

aH.J. Zhai and L.-S. Wang JACS 129 3022, 2007.

bS.Brnken; H. S. P. Mller; K.M. Menten; M. C. McCarthy and P. Thaddeus Ap. J. 676 1367, 2008.

cH.Wang; T.C. Steimle; C. Apetrei and J.P. Maier PCCP 11 2649, 2008.